# Progressive Tempering Sampler with Diffusion

### v1

@@ 1 Introduction
$10^{6}$$10^{7}$$10^{8}$$10^{9}$$10^{10}$$10^{11}$444455556666777788889999iDEMBNEMDDSCMCDPT+DM (baseline)PTSD (ours)Improved error and eval. $\leftarrow$ Target evaluation $\leftarrow$ Sample error ($\mathcal{W}_{2}$)
@@ 2 Background
$\,\mathrm{d}x_{t}{=}{-}2\dot{\sigma}(t)\sigma(t)\nabla_{x_{t}}\log p_{t}(x_{t})\,\mathrm{d}t{+}\sqrt{2\dot{\sigma}(t)\sigma(t)}\,\mathrm{d}\bar{w}_{t},$ (4)
@@ 3.1 Temperature guidance
$\vspace{-4pt}\frac{\partial}{\partial T}\nabla_{x_{t}}\log p_{t}(x_{t},T)\Big{|}_{T=T_{1}}\\ \approx\frac{\nabla_{x_{t}}\log p_{t}(x_{t},T_{2})-\nabla_{x_{t}}\log p_{t}(x_{t},T_{1})}{T_{2}-T_{1}}.$ (10)
@@ 3.2 Progressive Tempering Sampler with Diffusion
We now explore the use of temperature guidance to construct a pipeline that integrates neural samplers with parallel tempering (PT). Analogous to PT, we define a decreasing sequence of temperatures $[T_{K},T_{K-1},\cdots,T_{1}]$, where $T_{1}$ corresponds to the temperature of our target distribution.
@@ 3.3 Improving Techniques for Training PTSD
0:  Target density $\tilde{p}$, Temperatures $\{T_{k}\}_{k=1}^{K}$, Empty Buffers =$\mathcal{B}=\{\mathcal{B}_{k}\}_{k=1}^{K}$, Initial parallel tempering (PT) steps $L$, Refinement PT steps $l$, Truncate quantile $\tau$, Training iterations $M$, Buffer size $B$.
@@ 5.1 Extrapolation with Temperature guidance
The extrapolation is closer to the target temperature than either model at temperature $2.0$ or $1.5$, showing a large overlap with the ground truth. We also compare our exploration approach with several other strategies. (1) model extrapolation (ME): as we train a single model on both temperatures 2.0 and 1.5, and by conditioning on the temperature label, we can directly input temperature 1.0 and rely on the generalisation in deep neural networks, an approach that was demonstrated useful for interpolation in Moqvist et al. (2025). (2) auto-diff extrapolation (AD): we directly use Equation 9 instead of finite difference approximation for the time-derivative; (3) score rescaling heuristic (RS): we simply anneal the score with the temperature $\nabla_{x_{t}}\log p_{t}(x_{t},1.0)\approx 1.5\cdot\nabla_{x_{t}}\log p_{t}(x_{t},1.5)$, an approach considered by Skreta et al. (2025). Our proposed temperature guidance provides a better extrapolation than the other choices. Even though the extrapolation is not perfect, it serves as an initialization for the local PT refinement, as proposed in Sec. 3.3.
@@
FAB $0.23$±$\,\pm 0.01$ $0.30$±$\,\pm 0.02$ $2.50$±$\,\pm 0.19$ $0.32$±$\,\pm 0.01$ $0.16$±$\,\pm 0.02$ $5.70$±$\,\pm 0.01$
CMCD $0.14$±$\,\pm 0.01$ $0.08$±$\,\pm 0.01$ $3.36$±$\,\pm 0.22$ $0.69$±$\,\pm 0.01$ $0.62$±$\,\pm 0.01$ $7.44$±$\,\pm 0.03$
DDS $0.22$±$\,\pm 0.01$ $0.10$±$\,\pm 0.03$ $4.52$±$\,\pm 0.30$ $0.77$±$\,\pm 0.00$ $0.70$±$\,\pm 0.02$ $7.60$±$\,\pm 0.02$
iDEM 0.09±$\,\pm 0.01$ $\,\pm 0.01$ $3.26$±$\,\pm 0.42$ $0.87$±$\,\pm 0.01$ $0.87$±$\,\pm 0.01$ $8.30$±$\,\pm 0.02$ $0.49$±$\,\pm 0.01$ $1.95$±$\,\pm 0.00$
BNEM $0.12$±$\,\pm 0.00$ $0.07$±$\,\pm 0.00$ 2.16±$\,\pm 0.16$ $0.31$±$\,\pm 0.01$ $0.05$±$\,\pm 0.01$ $8.41$±$\,\pm 0.09$ 0.18±$\,\pm 0.01$ 0.01±$\,\pm 0.00$ 1.76±$\,\pm 0.00$
DiKL $\,\pm 0.01$ 0.030.03{0.03}0.03±$\,\pm 0.01$ $3.23$±$\,\pm 0.24$ $0.34$±$\,\pm 0.00$ $0.16$±$\,\pm 0.02$ $6.59$±$\,\pm 0.03$
PT+DM $0.13$±$\,\pm 0.05$ $\,\pm 0.06$ $3.01$±$\,\pm 0.66$ 0.13±$\,\pm 0.03$ 0.03±$\,\pm 0.02$ $\underline{5.04}$±$\,\pm 0.02$ 0.64±$\,\pm 0.02$ $0.22$±$\,\pm 0.01$ $1.82$±$\,\pm 0.00$
PTSD 0.08±$\,\pm 0.02$ 0.02±$\,\pm 0.01$ 1.93±$\,\pm 0.41$ 0.14±$\,\pm 0.05$ 0.04±$\,\pm 0.03$ 4.99±$\,\pm 0.07$ $0.79$±$\,\pm 0.01$ 0.19±$\,\pm 0.00$ 1.81±$\,\pm 0.00$
@@ 5.4 Scaling PTSD on Alanine Dipeptide
We also apply the method to the alanine dipeptide molecule in Cartesian coordinates. To maximise the benefits of PTSD, we run local PT refinement only for a subset of IS outputs, as detailed in Sec. 3.3. We evaluate both PTSD and PTDM with approximately the same amount of energy function evaluations ($2.6\times 10^{7}$), and show the resulting Ramachandran plots in Fig. 7. This example demonstrates the potential scalability and efficiency of PTSD in realistic sampling problems. We also evaluate the log-likelihood of ground-truth data under our PTDM and PTSD models, and the KL divergence between the ground-truth Ramachandran histogram and the model-generated one in Table 3. We show the projection of Ramachandran plots along two axes ($\phi$ and $\psi$) in Sec. D.4. As we can see, while PT+DM appears slightly better performance on generate the large metastable sates, resulting to be more plausible than PTSD along the $\phi$ axis, PTSD achieves better log-likelihood than PT+DM and also captures the small metastable state more effectively, as shown in Fig. 7. This showcases the great potential of PTSD when applied to more complex systems.
@@
Table 4: Ablations on the temperature guidance in Sec. 3.1 and the truncated IS in Sec. 3.3. We do not provide ablation on the local PT refinement as running MCMC is clearly helpful.
@@ Appendix A Target Distributions
is a mixture of Gaussians with 40 components in 2-dimensional space, proposed by Midgley et al. (2023). Each component of the MoG-40 has a mean within $[-40,40]\times[-40,40]$ and an diagonal covariance of $\texttt{softplus}(1)$.
@@ C.2 Hyperparameters
We evaluate iDEM using the implementation by Akhound-Sadegh et al. (2024). For all tasks, we use Eyler-Maruyama discretization with $1000$ steps and $100$ inner-loops. For MoG-40, we train iDEM for $1000$ outer-loops to ensure convergence, with a batch size of $1000$. The marginal scores are estimated by $500$ MC samples, which is clipped to a maximum norm of $70$. The total number of target evaluations is $1000\times 100\times 1000\times 5000$. For MW-32, we train iDEM for $180$ outer-loops, increase the number of MC samples to $1000$, clip the score to a maximum norm of $1000$, and keep the other hyperparameter the same. The total number of target evaluations is $180\times 100\times 1000\times 1000$. For LJ-55, we further clip the score to a maximum norm of $20$, decrease the batch size to $128$, and train iDEM for $1000$ outer-loops. The total number of target evaluations is $1000\times 100\times 128\times 1000$.
@@ D.1 Ablation Study for Extrapolation
We conduct experiments on GMM-40, DW-4  [^4] DW-4 is a 4-particle system in 2-dimensional space introduced by Köhler et al. (2020). To clarify, we follow the notation used in (Köhler et al., 2020; Akhound-Sadegh et al., 2024) where the number $4$ indicates the number of particles. While the number $2$ in DW-2 indicates the potential is in a 2-dimensional space., LJ-13, and LJ-55 benchmarks. For all experiments, we choose $T_{1}=0.1$, $T_{2}=1$, and $T_{3}=1.3$, where $T_{2}=1$ is the same as the target distributions introduced in App. A. We use the same architecture introduced in Sec. C.1, and the training of each model is long enough to ensure convergence. The ground truth of each benchmark at $T_{1}=0.1$ is obtained by running MCMC on top of the $T_{2}=1$ samples for a long enough time to ensure mixing.
@@ D.2 Measuring Performance through Observables
FAB 1.91±$\,\pm 0.98$ $4.09$±$\,\pm 0.14$
CMCD $7.03$±$\,\pm 0.83$ $3.76$±$\,\pm 0.12$
DDS $6.85$±$\,\pm 1.36$ $9.89$±$\,\pm 0.13$
iDEM $6.5$±$\,\pm 0.75$ $9.28$±$\,\pm 0.06$ $28.72$±$\,\pm 0.02$
BNEM $7.42$±$\,\pm 0.63$ $11.09$±$\,\pm 0.18$ $9.48$±$\,\pm 0.01$
DiKL $7.59$±$\,\pm 0.76$ 3.24±$\,\pm 0.17$
PT+DM $12.09$±$\,\pm 0.81$ 4.31±$\,\pm 0.10$ 0.34±$\,\pm 0.03$
PTSD (ours) 1.35±$\,\pm 0.69$ 1.49±$\,\pm 0.16$ 1.86±$\,\pm 0.04$
@@ D.3 Measuring Performance through Log-Likelihood
[…]

### v2

@@ 1 Introduction
[…]
@@ 2 Background
$\hskip-8.99994pt\,\mathrm{d}x_{t}{=}{-}2\dot{\sigma}(t)\sigma(t)\nabla_{x_{t}}\log p_{t}(x_{t})\,\mathrm{d}t{+}\sqrt{2\dot{\sigma}(t)\sigma(t)}\,\mathrm{d}\bar{w}_{t},$ (4)
@@ 3.1 Temperature guidance
$\vskip-4.0pt\frac{\partial}{\partial T}\nabla_{x_{t}}\log p_{t}(x_{t},T)\Big|_{T=T_{1}}\\ \approx\frac{\nabla_{x_{t}}\log p_{t}(x_{t},T_{2})-\nabla_{x_{t}}\log p_{t}(x_{t},T_{1})}{T_{2}-T_{1}}.$ (10)
@@ 3.2 Progressive Tempering Sampler with Diffusion
We now explore the use of temperature guidance to construct a pipeline that integrates neural samplers with parallel tempering (PT). Analogous to PT, we define a decreasing sequence of temperatures $T_{K},T_{K-1},\cdots,T_{1}]$, where $T_{1}$ corresponds to the temperature of our target distribution.
@@ 3.3 Improving Techniques for Training PTSD
0: Target density $\tilde{p}$, Temperatures $\{T_{k}\}_{k=1}^{K}$, Empty Buffers $\mathcal{B}=\{\mathcal{B}_{k}\}_{k=1}^{K}$, Initial parallel tempering (PT) steps $L$, Refinement PT steps $l$, Truncate quantile $\tau$, Training iterations $M$, Buffer size $B$.
@@ 5.1 Extrapolation with Temperature guidance
The extrapolation is closer to the target temperature than either model at temperature $2.0$ or $1.5$, showing a large overlap with the ground truth. We also compare our exploration approach with several other strategies. (1) model extrapolation (ME): as we train a single model on both temperatures 2.0 and 1.5, and by conditioning on the temperature label, we can directly input temperature 1.0 and rely on the generalisation in deep neural networks, an approach that was demonstrated useful for interpolation in Moqvist et al. (2025). (2) auto-diff extrapolation (AD): we directly use Equation 9 instead of finite difference approximation for the time-derivative; (3) score rescaling heuristic (RS): we simply anneal the score with the temperature $\nabla_{x_{t}}\log p_{t}(x_{t},1.0)\approx 1.5\cdot\nabla_{x_{t}}\log p_{t}(x_{t},1.5)$, an approach considered by Skreta et al. (2025). Our proposed temperature guidance provides a better extrapolation than the other choices. Even though the extrapolation is not perfect, it serves as an initialization for the local PT refinement, as proposed in Sec. $3.$
@@
[…]
@@ 5.4 Scaling PTSD on Alanine Dipeptide
We also apply the method to the alanine dipeptide molecule in Cartesian coordinates. To maximise the benefits of PTSD, we run local PT refinement only for a subset of IS outputs, as detailed in Sec. $3.$ We evaluate both PTSD and PTDM with approximately the same amount of energy function evaluations ($2.6\times 10^{7}$), and show the resulting Ramachandran plots in Fig. 7. This example demonstrates the potential scalability and efficiency of PTSD in realistic sampling problems. We also evaluate the log-likelihood of ground-truth data under our PTDM and PTSD models, and the KL divergence between the ground-truth Ramachandran histogram and the model-generated one in Table 3. We show the projection of Ramachandran plots along two axes ($\phi$ and $\psi$) in Sec. D.4. As we can see, while PT+DM appears slightly better performance on generate the large metastable sates, resulting to be more plausible than PTSD along the $\phi$ axis, PTSD achieves better log-likelihood than PT+DM and also captures the small metastable state more effectively, as shown in Fig. 7. This showcases the great potential of PTSD when applied to more complex systems.
@@
Table 4: Ablations on the temperature guidance in Sec. 3.1 and the truncated IS in Sec. $3.$ We do not provide ablation on the local PT refinement as running MCMC is clearly helpful.
@@ Appendix A Target Distributions
is a mixture of Gaussians with 40 components in 2-dimensional space, proposed by Midgley et al. (2023). Each component of the MoG-40 has a mean within $40,40]\times[-40,40]$ and an diagonal covariance of $\texttt{softplus}(1)$.
@@ C.2 Hyperparameters
We evaluate iDEM using the implementation by Akhound-Sadegh et al. (2024). For all tasks, we use Eyler-Maruyama discretization with $1000$ steps and $100$ inner-loops. For MoG-40, we train iDEM for $1000$ outer-loops to ensure convergence, with a batch size of $1000$. The marginal scores are estimated by $500$ MC samples, which is clipped to a maximum norm of $70$. The total number of target evaluations is $1000\times 100\times 1000\times 5000$. For MW-32, we train iDEM for $180$ outer-loops, increase the number of MC samples to $1000$, clip the score to a maximum norm of $1000$, and keep the other hyperparameter the same. The total number of target evaluations is $180\times 100\times 1000\times 1000$. For LJ-55, we further clip the score to a maximum norm of 2020, decrease the batch size to $128$, and train iDEM for $1000$ outer-loops. The total number of target evaluations is $1000\times 100\times 128\times 1000$.
@@ D.1 Ablation Study for Extrapolation
We conduct experiments on GMM-40, DW-4  [^4] DW-4 is a 4-particle system in 2-dimensional space introduced by Köhler et al. (2020). To clarify, we follow the notation used in (Köhler et al., 2020; Akhound-Sadegh et al., 2024) where the number 44 indicates the number of particles. While the number 22 in DW-2 indicates the potential is in a 2-dimensional space., LJ-13, and LJ-55 benchmarks. For all experiments, we choose $T_{1}=0.1$, $T_{2}=1$, and $T_{3}=1.3$, where $T_{2}=1$ is the same as the target distributions introduced in App. A. We use the same architecture introduced in Sec. C.1, and the training of each model is long enough to ensure convergence. The ground truth of each benchmark at $T_{1}=0.1$ is obtained by running MCMC on top of the $T_{2}=1$ samples for a long enough time to ensure mixing.
@@ D.2 Measuring Performance through Observables
[…]
@@ D.3 Measuring Performance through Log-Likelihood
$\log p_{\theta}(x_{0})$ $=\log p_{1}(x_{1})+\int_{0}^{1}\nabla\cdot\tilde{f}(x_{t},t)dt,\quad\text{with }x_{t}=x_{1}-\int_{1}^{t}\tilde{f}(x_{u},u)du.$ (20)